# Periodic orbits and their gravitational wave radiations in black hole with dark matter halo

### v1

@@ V Gravitational Waveform from Periodic Orbits
EMRIs are promising sources for future space-based GW detectors. These systems, comprising a stellar-mass compact object orbiting an SMBH, emit GWs that encode rich information about the system’s dynamics and the surrounding spacetime. When the smaller object is in a periodic orbit around an SMBH immersed in a DM halo, the emitted GW waveform offers a unique opportunity to study distinct features that can be exploited to probe the system’s properties. This section outlines the theoretical framework for calculating gravitational waveforms from such periodic orbits.
The computation of gravitational waveforms from EMRIs often relies on the adiabatic approximation, which is valid when the inspiral timescale is much longer than the orbital period. Under this approximation, the smaller object’s energy and angular momentum can be considered constant over several orbits, allowing the trajectory to be modeled as a geodesic in the background spacetime of the SMBH. The back reaction of gravitational radiation on the motion of the smaller object is effectively neglected Hughes:1999bq ; Isoyama:2021jjd .
Figure 9: Gravitational waveforms from a test object with $m=10M_{\odot}$ around periodic orbits (1,2,0): blue, (2,1,1): red, (3,2,2): orange, around an SMBH immersed in a DM halo with $M_{B}=10^{6}M_{\odot}$. Other parameters are $M_{H}=10$ and $a_{0}=1000$. The energy is fixed at $E=0.96$.
A numerical “kludge” waveform model provides a practical approach to calculating the GWs from these periodic orbits within the DM halo environment Babak:2006uv . This method involves a two-step process: first, the orbit of the small object is determined by numerically solving the geodesic equations of motion in the black hole’s spacetime, accounting for the DM distribution. Subsequently, the quadrupole formula for gravitational radiation is applied to this calculated orbit, generating the corresponding waveform. This approach allows for a preliminary investigation of the GW signals emitted by EMRIs and their potential to reveal insights into the properties of both the orbit and the central black hole, as well as the distribution of DM surrounding it.
@@
where $A=c^{4}D_{L}/(2G)$, $G=1=c$, $D_{L}$ is the luminosity distance to the source. By numerically solving the geodesic equations of motion, the trajectory $Z_{i}(t)$ of the smaller object in the curved spacetime of the SMBH immersed in DM halo can be determined. This trajectory is crucial for the subsequent calculation of the gravitational waveform. For a small object of mass $m$ following a trajectory $Z^{i}(t)$, the $I_{ij}$ is given by Thorne:1980ru
@@
The choice of coordinate system plays a crucial role in the calculation and interpretation of gravitational waveforms. While the geodesic equations are often solved in Boyer-Lindquist coordinates $(r,\theta,\phi)$, the waveform is typically expressed in a detector-adapted coordinate system $(X,Y,Z)$. This transformation simplifies the analysis of the signal as measured by a GW detector. The transformation from Boyer-Lindquist to Cartesian coordinates is given by Babak:2006uv ,
@@
VI Conclusion
@@
Finally, we analyze an EMRI system consisting of a test object with mass $m=10M_{\odot}$ following periodic orbits around a SMBH, having mass $M_{B}=10^{6}M_{\odot}$, immersed in DM halo. Using the numerical kludge scheme, we investigated the resulting gravitational waveforms by positioning the system at a luminosity distance of $D_{L}=2$ Gpc from the detector, with an inclination angle of $\iota=\pi/4$ and a longitude of pericenter $\zeta=\pi/4$. This study demonstrates a clear correlation between the gravitational waveforms emitted by a small object orbiting a SMBH and the object’s zoom-whirl orbital behavior. Higher zoom numbers correspond to more complex waveform substructures. Furthermore, the presence of a DM halo significantly impacts these waveforms. Decreasing the halo’s scale length $a_{0}$ or increasing its mass $M_{H}$ results in both a phase shift and amplitude increase in the waveforms.

### v2

@@ V Gravitational Waveform from Periodic Orbits
EMRIs are promising sources for future space-based GW detectors. These systems, comprising a stellar-mass compact object orbiting a SMBH, emit GWs that encode rich information about the system’s dynamics and the surrounding spacetime. When the smaller object is in a periodic orbit around a SMBH immersed in a DM halo, the emitted GW waveform offers a unique opportunity to study distinct features that can be exploited to probe the system’s properties. This section outlines the theoretical framework for calculating gravitational waveforms from such periodic orbits.
The computation of gravitational waveforms from EMRIs often relies on the adiabatic approximation, which is valid when the inspiral timescale is much longer than the orbital period. Under this approximation, the smaller object’s energy and angular momentum can be considered constant over several orbits, allowing the trajectory to be modeled as a geodesic in the background spacetime of the SMBH. The back reaction of gravitational radiation on the motion of the smaller object is effectively neglected [76, 77].
Figure 9: Gravitational waveforms from a test object with $m=10M_{\odot}$ around periodic orbits (1,2,0): blue, (2,1,1): red, (3,2,2): orange, around a SMBH immersed in a DM halo with $M_{B}=10^{6}M_{\odot}$. Other parameters are $M_{H}=10$ and $a_{0}=1000$. The energy is fixed at $E=0.96$.
A numerical “kludge” waveform model provides a practical approach to calculating the GWs from these periodic orbits within the DM halo environment [75]. This method involves a two-step process: first, the orbit of the small object is determined by numerically solving the geodesic equations of motion in the black hole’s spacetime, accounting for the DM distribution. Subsequently, the quadrupole formula for gravitational radiation is applied to this calculated orbit, generating the corresponding waveform. This approach allows for a preliminary investigation of the GW signals emitted by EMRIs and their potential to reveal insights into the properties of both the orbit and the central black hole, as well as the distribution of DM surrounding it.
@@
where $A=c^{4}D_{L}/(2G)$, $G=1=c$, $D_{L}$ is the luminosity distance to the source. By numerically solving the geodesic equations of motion, the trajectory $Z_{i}(t)$ of the smaller object in the curved spacetime of the SMBH immersed in DM halo can be determined. This trajectory is crucial for the subsequent calculation of the gravitational waveform. For a small object of mass $m$ following a trajectory $Z^{i}(t)$, the $I_{ij}$ is given by [78]
@@
Figure 12: Absolute frequency spectra $|\tilde{h}_{+,\times(f)}|$ corresponding to the waveforms in Figure 9.
Figure 13: Absolute frequency spectra $|\tilde{h}_{+,\times(f)}|$ corresponding to the waveforms in Figure 10.
Figure 14: Absolute frequency spectra $|\tilde{h}_{+,\times(f)}|$ corresponding to the waveforms in Figure 11.
The choice of coordinate system plays a crucial role in the calculation and interpretation of gravitational waveforms. While the geodesic equations are often solved in Boyer-Lindquist coordinates $(r,\theta,\phi)$, the waveform is typically expressed in a detector-adapted coordinate system $(X,Y,Z)$. This transformation simplifies the analysis of the signal as measured by a GW detector. The transformation from Boyer-Lindquist to Cartesian coordinates is given by [75]
@@
[…]
$h_{c}(f)=2f\left(|\tilde{h}_{+}(f)|^{2}+|\tilde{h}_{\times}(f)|^{2}\right)^{1/2},$ (5.15)
To obtain the frequency spectra, we apply a discrete Fourier transform (DFT) to the time-domain gravitational waveforms, converting the signal from the time domain to the frequency domain. This allows for a detailed examination of the frequency distribution present in the signal, revealing the influence of the periodic orbital motion of the particle on the frequency structure of the gravitational wave signal, see Figs. 12, 13 and 14. The characteristic frequencies of gravitational waves from these periodic orbits fall primarily within the millihertz range, which is particularly significant for space-based gravitational wave detectors such as LISA, optimized for detecting low-frequency gravitational waves. The characteristic frequencies for different periodic orbits, indexed by $(z,w,v)$, are shown in Fig. 12, while the characteristic frequencies for various values of the parameters $a_{0}$ and $M_{H}$ are depicted in Figs. 13 and 14, respectively.
Figure 15: Characteristic strain $h_{c}(f)$ of gravitational waveforms of periodic orbits in Figures 9, 10 and $1$.
To enhance the visual clarity of the plots for characteristic strain (5.15), we applied a smoothing technique to the numerically generated $h_{c}(f)$ by performing a running average over 30 frequency bins. Selecting a larger window size would have smoothed out more of the noise but may have blurred some of the finer features in the data. It can be observed, in Fig. 15, that some segments of the characteristic strain, corresponding to different orbital configurations $(z,w,v)$ and parameters $a_{0}$ and $M_{H}$, exceed the sensitivity curve of LISA (Laser Interferometer Space Antenna). This suggests that these gravitational waves, characterized by their distinctive zoom-whirl features in the presence of a dark matter halo, are likely to be detectable by LISA, offering valuable insights into the spacetime geometry surrounding supermassive black holes.
VI Conclusion and Outlook
@@
Finally, we analyze an EMRI system consisting of a test object with mass $m=10M_{\odot}$ following periodic orbits around a SMBH, having mass $M_{B}=10^{6}M_{\odot}$, immersed in DM halo. Using the numerical kludge scheme, we investigated the resulting gravitational waveforms by positioning the system at a luminosity distance of $D_{L}=2$ Gpc from the detector, with an inclination angle of $\iota=\pi/4$ and a longitude of pericenter $\zeta=\pi/4$. This study demonstrates a clear correlation between the gravitational waveforms emitted by a small object orbiting a SMBH and the object’s zoom-whirl orbital behavior. Higher zoom numbers correspond to more complex waveform substructures. Furthermore, the presence of a DM halo significantly impacts these waveforms. Decreasing the halo’s scale length $a_{0}$ or increasing its mass $M_{H}$ results in both a phase shift and amplitude increase in the waveforms. To assess how well gravitational waves from EMRIs with periodic orbits can be detected, we analyzed their time-domain waveforms using discrete Fourier transforms to extract the frequency spectra. The results indicate that the frequencies of these gravitational waves generally fall within the sensitivity range of space-based detectors. From the spectra, we determined the characteristic strains and observed that, for certain combinations of $(z,w,v)$ or specific values of the parameters $a_{0}$ and $M$, the strains surpass LISA’s sensitivity threshold. This suggests that space-based gravitational wave observatories could capture signals from EMRIs with periodic orbits, providing a promising avenue to explore supermassive black holes with dark matter halos.
@@
Here we would like to make a few remarks about the limitation of the calculations of the waveforms and the potential extensions of the current study. First, we use the adiabatic approximation with which we ignore the backreaction of the gravitational radiation to the periodic orbits, which is a valid approximation when considering only a few orbital periods, as is the case in this study. Exploring the impact of gravitational radiation on the long-term evolution of periodic orbits represents an interesting avenue for future research. Second, we also ignores the contributions of multipoles higher than the quadratic order. It is also crucial for developing more accurate waveform models to include higher-order multipole moments in the gravitational wave expansion. At last, once these accurate waveforms become available, we will be able to investigate how future gravitational wave detectors might constrain or test the effects of dark matter on periodic orbits. We look forward to addressing these challenges in subsequent studies.